OpenStudy (anonymous):

Sides of three square rooms measure 14 feet each, and sides of two square rooms measure 17 feet each. Which expression shows the total area of these five rooms? A.)(3 × 14^2) + (2 × 17^2) B.)(2 × 14^3) + (2 × 17^2) C.)(3 × 17^2) + (2 × 14^2) D.)(3 × 14^2) × (2 × 17^2)

OpenStudy (anonymous):

@Learner11 im thinking b

OpenStudy (anonymous):

Can you explain your answer to me @magicremix123

OpenStudy (anonymous):

you are almost done

OpenStudy (anonymous):

wait...its (a) because the area of a square is the product of two of its sides the 13 foot square = 13*13 = 13^2 for three squares= 3* 13^2 the 15 foot square = 15*15 = 15^2 for 2 squares = 2 * 15^2

OpenStudy (anonymous):

right? i believe so, i recognized it afterwards..pretty easy, didn't realize

OpenStudy (anonymous):

YUp ! Well done :D . Your explanation is absolutely correct . I will give you a medal . But I think those are 14 and 17 instead of 13 and 15 . Isn't it ?

OpenStudy (anonymous):

Yes, but my teacher has me doing it both ways.

OpenStudy (anonymous):

Thank you, for taking your time to look over my answer. :)

OpenStudy (anonymous):

it's Completely fine :D